Select all of the following solutions that would form a buffer when they were mixed.

A solution that is formed when 0.82 moles of NH3 is mixed with 0.24 moles of HBr. Assume the final volume is 1.0 L.

A solution that is formed when 0.52 moles of NaNO2 is mixed with 0.24 moles of HCl. Assume the final volume is 1.0 L.

A solution that is formed when 0.23 moles of CH3CH2COOH is mixed with 0.49 moles of NaOH. Assume the final volume is 1.0 L.

A solution that is formed when 0.33 moles of NH3 is mixed with 0.12 moles of NaOH. Assume the final volume is 1.0 L.

A solution that is formed when 0.33 moles of CH3COOH is mixed with 0.39 moles of NaOH. Assume the final volume is 1.0 L.
Answer #1

A solution that is formed when 0.82 moles of NH3 is mixed with 0.24 moles of HBr. Assume the final volume is 1.0 L. (buffer)

mole of NH4Br = 0.24 mole and mole of NH3 = (0.82 - 0.24) = 0.58 mole.

so this solution is buffer solution.

A solution that is formed when 0.52 moles of NaNO2 is mixed with 0.24 moles of HCl. Assume the final volume is 1.0 L. (buffer)

In this solution, mole of HNO2 = 0.24 mole and mole of NaNO2 = (0.52 - 0.24) = 0.28 mole.

this is buffer solution.

A solution that is formed when 0.23 moles of CH3CH2COOH is mixed with 0.49 moles of NaOH. Assume the final volume is 1.0 L.

there are excess NaOH. so this is not buffer solution.

A solution that is formed when 0.33 moles of NH3 is mixed with 0.12 moles of NaOH. Assume the final volume is 1.0 L.

this is not buffer as there both NH3 and NaOH are base.

A solution that is formed when 0.33 moles of CH3COOH is mixed with 0.39 moles of NaOH. Assume the final volume is 1.0 L.

there are excess NaOH. so this is not buffer solution.